OpenStudy (anonymous):

Joan, who is 15 years old, receives an inheritance of $21,000 from her aunt Sally. Her parents help her to invest the money in a fund which doubles her money every 12 years. What age will Joan be when the account has grown to $500,000?

OpenStudy (tkhunny):

Easy Way: \(500000 = 21000(2^{n/12})\). -- Solve for n. Add 15. More Standard Way: \(2*21000 = 21000\cdot e^{12k}\) -- Solve for k. \(500000 = 21000\cdot e^{nk}\) -- Solve for n. Add 15.

OpenStudy (anonymous):

Will she be about 30?

OpenStudy (tkhunny):

You tell me. How did you get that? If it doubles in 12, 21000 ==> 42000 at age 27, it seems unlikely that it will double several more times in the next 3 years. Make sure your answer makes sense. This will save you from marking errors.

OpenStudy (anonymous):

Oh okay so about 70 yrs old then

OpenStudy (anonymous):

Wait no.... anout 55 right. Yeah its 55 yrs old

OpenStudy (anonymous):

Nope its 70. 55 doesn't make sense

OpenStudy (tkhunny):

69.8816 Again, how did you do that? What did you do? What were your intermediate results? You have to tell us what is going on. We can't see in your head.

OpenStudy (anonymous):

I got around 70 because I doubled the money and added 12 years. Thanks

OpenStudy (tkhunny):

21000*2 = 42000 -- 12 years age 27 42000*2 = 84000 -- 12 years age 39 84000*2 = 178000 -- 12 years age 51 178000*2 = 356000 -- 12 years age 63 356000*2 = 712000 -- 12 years age 75 -- Whoops! Too far. I DO wish I could tell it was what you were thinking. Unfortunately, you will not spend enough time telling me. Good luck.